A gas has a volume of 1.4 liters at STP. What is the new volume when the temperature is changed to 40 K and the pressure is changed to 1.2 atm?
1
Expert's answer
2020-04-28T12:20:44-0400

STP conditions are:

"P_1=1\ atm,V_1=22.4\ L,T_1=298\ K"

New conditions are:

"P_2=1.2 \ atm,V_2=?,T_2=40\ K"

Using ideal gas equation,

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}\implies \frac{1\times 22.4}{298}=\frac{1.2\times V_2}{40}" "\implies V_2=2.50\ L"